OpenStudy (anonymous):

a gym has a special offer on enrollment. there i a one time registration fee of 75 dollars and the you pay 38 dollars per month. Although,Dr. Jamal has 500 dollars.what is the maximum number of months he can work out at the gym?

OpenStudy (anonymous):

75 + 38m <= 500 --- m stands for # of months There is your inequality...can you solve it from here ?

OpenStudy (anonymous):

ill will try

OpenStudy (anonymous):

how ya doing ? Need any help ?

OpenStudy (anonymous):

what did you get

OpenStudy (anonymous):

im stuck

OpenStudy (anonymous):

11 months...what about you ? I did this : 75 + 38m <= 500 38m <= 500 - 75 38m <= 425 m <= 425/38 m <= 11.18.....round down to 11 so it has to be 11 months OR LESS \any questions ? any at all ?

OpenStudy (anonymous):

oh....the maximum...it would be 11 months maximum

OpenStudy (anonymous):

check.. 75 + 38(11) <= 500 75 + 418 <= 500 493 <= 500 as you can see, if he goes 11 months, that will leave him with $7 less then $500

OpenStudy (anonymous):

i thought it would be 14 months